Circular No. 6031

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) For the Hubble Space Telescope observing team (HST), K. S. Noll,

Space Telescope Science Institute, reports the following results: “Ultraviolet spectra of the central 0′′.9 of impact site G obtained with the HST Faint Object Spectrograph (FOS) on July 18 show a series of strong absorption bands over 255–293 nm that we unambiguously identify as due to S_{2}. A second strong absorption occurs between 180 and 230 nm caused by a combination of enhanced NH_{3} and at least one other molecule. The absorption cross section of H_{2}S provides a good match to the spectrum, and we regard this as a tentative identification. The absorption bands are observable in both I/F and in ratios of the impact site spectrum to the spectrum of the same location obtained on July 14. The spectral resolution of the FOS for an extended source observed through the 0′′.9 aperture is 0.4 nm over 160–230 nm and 0.6 nm over 230–330 nm. We note that enhancements of NH_{3}, H_{2}S and S_{2} could result from heating and vaporization of the NH_{4}SH and NH_{3} clouds. If the sulphur were cometary in origin, we should also expect enhanced abundances of oxygen- and carbon-bearing compounds to be detected.”

No report has been received of the impact on Jupiter of fragment P = 8, but one received via the SL9 message center notes that the French-Spanish-Swedish observing team recorded the impact of Q = 7 (i.e., Q_{1} = 7a) at 10 μm with the Saclay CAMIRAS camera mounted on the Nordic Optical Telescope at La Palma. The event occurred on July 20.848 UT and was less bright than the L impact seen yesterday. M. Lindgren, Uppsala Astronomical Observatory, reports, also from La Palma, that the Q impact site was in view on July 20.854 UT. V band imaging with the Swedish solar telescope shows it to be very similar to the nearby L site: a central dark spot and a diffuse concentric ring. The seeing was not very good, but it was just about possible to see an elongation of the central dark spot, perhaps due to the double impact of Q_{1} and Q_{2}.

T. Z. Martin has reported via the SL9 message center the attempted detection of the impacts of fragments B, H and L by the photopolarimeter radiometer on the Galileo spacecraft (using a filter at 945 nm). No evidence was found of the B impact. The H and L impacts were observed at equivalent earth times July 18.81388 and 19.92833 UT, the signal rising to a peak value in ∼ 2 s and then decreasing to background levels over 25 and 35 s, respectively. Peak brightness corresponded to ∼ 2 and 4 percent, respectively, of the brightness of Jupiter itself.

1994 July 20 Brian G. Marsden